Integrate dx/(4+x²)½ using trigonometric substitution
1
Expert's answer
2020-02-24T12:22:13-0500

Integrate d x/(4+x²)½ using trigonometric substitution

Suppose "tan\theta={x \above{2pt} 2}"

by applying cross multiplication

x=2"tan\theta"

"dx=2sec^2\theta"

"\int {dx\above{2pt}\sqrt{x^2+4} }"

"=\int{2sec^2 \theta\above{2pt}\sqrt{4tan^2\theta+4} } d\theta"

"=\int{2sec^2\theta \above{2pt}\sqrt{4(tan^2\theta+1)} } d\theta"

using trigonometric identity "tan^2\theta+1=sec^2\theta"

"=" "\int{2sec^2\theta\above{2pt}\sqrt{4(sec^2\theta)} }d\theta"

"=\int{2sec^2\theta\above{2pt} 2sec\theta} d\theta"

"=" "\int sec\theta d\theta"

multiplying and divided both sides by "sec\theta+tan\theta"

"=\int sec\theta({sec\theta+tan\theta \above{2pt} sec\theta+tan\theta}) d\theta"

"=\int{ sec^2\theta+tan\theta sec\theta\above{2pt} sec\theta+tan\theta}d\theta"

Substitute u="sec\theta+tan\theta" (A)

"du=(sec\theta tan\theta+sec^2\theta)d\theta"

"=\int {du \above{2pt} u}"

"=ln|u|+C"

By putting the value of u

"=ln|sec\theta+tan\theta|+C" (1)


Replace "\theta" terms with x

"tan\theta" "={x \above{2pt} 2}"

taking square on both sides

"tan^2\theta=({x \above{2pt} 2})^2"

"sec^2\theta-1={x^2 \above{2pt} 4}"

"sec^2\theta={x^4\above{2pt} 4}+1"

"sec^2\theta={x^2+4 \above{2pt} 4}"

taking square root on both sides

"sec\theta={ \sqrt{x^2+4}\above{2pt} 2}"

By putting the value in equation no 1

"=ln( { \sqrt{x^2+4} \above{2pt} 2}+{x\above{2pt} 2} )+C" Answer